# Instance Temperature Knowledge Distillation

### v3

@@ Abstract
Knowledge distillation (KD) enhances the performance of a student network by allowing it to learn the knowledge transferred from a teacher network incrementally. Existing methods dynamically adjust the temperature to enable the student network to adapt to the varying learning difficulties at different learning stages of KD. KD is a continuous process, but when adjusting the temperature, these methods consider only the immediate benefits of the operation in the current learning phase and fail to take into account its future returns. To address this issue, we formulate the adjustment of temperature as a sequential decision-making task and propose a method based on reinforcement learning, termed RLKD. Importantly, we design a novel state representation to enable the agent to make more informed action (i.e., instance temperature adjustment). To handle the problem of delayed rewards in our method due to the KD setting, we explore an instance reward calibration approach. In addition, we devise an efficient exploration strategy that enables the agent to learn valuable instance temperature adjustment policy more efficiently. Our framework can serve as a plug-and-play technique to be inserted into various KD methods easily, and we validate its effectiveness on both image classification and object detection tasks. Our project is at https://www.zayx.me/ITKD.github.io/.
@@ 1 Introduction
Over the past few decades, the field of computer vision has undergone a transformative shift thanks to the remarkable progress of deep neural networks (DNNs). Nonetheless, the significant computation and storage demands of DNNs [26, 45, 33, 34], present great challenges, especially in industrial applications where there is a preference for efficient and lightweight models. Typically, lightweight networks do not perform as well as deeper networks. To solve this issue, knowledge distillation (KD), which aims to enable smaller (student) models to compete with larger (teacher) models in performance, has been introduced [14]. Due to its remarkable effectiveness in boosting the capability of lightweight models, KD has been widely used in various tasks, e.g., object detection [7, 42], semantic segmentation [24, 39, 44], and natural language processing [31, 11].
@@
In summary, our main contributions are: 1) In KD, to account for the future benefits of adjusting instance temperature at the current stage, we formulate the instance temperature adjustment as a sequential decision-making task, and propose a novel method RLKD based on RL to handle this task. 2) To overcome the challenge of delayed rewards in our RLKD, we exploit a mechanism for instance reward calibration. Furthermore, we design a valid exploration strategy to promote the agent to learn valuable temperature adjustment policy with high efficiency. 3) Our RLKD can serve as a plug-and-play technique to boost the performance of KD algorithms. We validate its effectiveness on three benchmarks for image classification and object detection, all obtaining the state-of-the-art result.
@@ 3 Preliminary on Reinforcement Learning
$\vspace{-0.7mm}G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\vspace{-0.4mm}$ (1)
@@ 4.2 Instance Reward Calibration
Here, $r^{b^{\prime}}_{n}$ represents the $n$-th corrected reward, and $n$ is the batch size. The variables $alpha$ and $beta$ are weights, which we set to 1 and 0.5, respectively. The variable $G_{i}$ denotes the return at the $i$-th time step. Additionally, to ensure that the states $s^{b}$ recorded in the replay buffer match the corrected rewards, we devise a state updater $\mathcal{U}$ and update the states $s^{b}$ accordingly. The updated state $s^{b^{\prime}}$ is calculated as follows:
@@ 4.3 Efficient Exploration
Input: agent network $Q$, student network $f_{S}$, teacher network $f_{T}$, dataset $D$, batch size $N$, reward corrector $\mathcal{C}$
for batch $d_{t+1}$ in $D$ do
       for $i=0$ to $N-1$ do
             Build state observation $s_{i}$ from $f_{S}(d_{t+1}^{i})$, $f_{T}(d_{t+1}^{i})$
             Compute the action $a_{i}$ and values $V_{i}\leftarrow Q(s_{i})$
             Use $a_{i}$ as instance temperature
             Collect $s_{i}$, $a_{i}$, $V_{i}$ as a buffer to update $Q$ (Sec. 4.1)
       end for
      Update the student network $f_{S}$
       Obtain the reward $r_{t+1}$ following Eq. 9
       Calibrate the instance reward following Eq. 10
       Update the reward $r_{t+1}$ in the replay buffer
       Update the state $s_{t}$ following Eq. 12 (Sec. 4.2)
       while not done do
             Update $Q$
       end while
end for
Compute prediction entropy $H_{t}\leftarrow f_{S}(d_{t+1})$
Obtain high-quality samples following Eq. 16 (Sec. 4.3)
Implement the efficient exploration strategy
Update $Q$
@@ 5 Experiments
Tasks and datasets. Following [20], we conduct experiments on two tasks: image classification and object detection. For the image classification task, we carry out experiments on CIFAR-100 [18] and ImageNet [8]. For the object detection task, we conduct our experiments on MS-COCO [21]. CIFAR-100 is a prominent dataset for image classification, comprising 32$\times$32 pixel images across 100 different categories, with a training set of 50,000 images and a validation set of 10,000 images. ImageNet, another significant dataset for large-scale image classification, encompasses 1,000 categories with a training set of approximately 1.28 million images and a validation set of 50,000 images. MS-COCO is a famous dataset used for general object detection that includes 80 categories. It has a training set (train 2017) with 118,000 images and a validation set (val 2017) with 5,000 images.
@@ 5.1 Main results
Table 7: Ablation study of the efficient exploration (EE) strategy.
@@ 5.2 Ablation studies
Table 8: Comparison of different high-quality training sample selection strategies. The teacher-student pairs corresponding to the second and third rows are respectively RN-56 & RN-20 and WRN-40-2 & WRN-16-2. “$\mathcal{M}$” denotes the mix-up operation.

### v5

@@ Abstract
Knowledge distillation (KD) enhances the performance of a student network by allowing it to learn the knowledge transferred from a teacher network incrementally. Existing methods dynamically adjust the temperature to enable the student network to adapt to the varying learning difficulties at different learning stages of KD. KD is a continuous process, but when adjusting the temperature, these methods consider only the immediate benefits of the operation in the current learning phase and fail to take into account its future returns. To address this issue, we formulate the adjustment of temperature as a sequential decision-making task and propose a method based on reinforcement learning, termed RLKD. Importantly, we design a novel state representation to enable the agent to make more informed action (i.e., instance temperature adjustment). To handle the problem of delayed rewards in our method due to the KD setting, we explore an instance reward calibration approach. In addition, we devise an efficient exploration strategy that enables the agent to learn valuable instance temperature adjustment policy more efficiently. Our framework can serve as a plug-and-play technique to be inserted into various KD methods easily, and we validate its effectiveness on both image classification and object detection tasks. Our code is at https://itkd123.github.io/ITKD.github.io/.
 [^1] First Author and Second Author contribute equally to this work.
@@ 1 Introduction
Over the past few decades, the field of computer vision has undergone a transformative shift thanks to the remarkable progress of deep neural networks (DNNs). Nonetheless, the significant computation and storage demands of DNNs, present great challenges, especially in industrial applications where there is a preference for efficient and lightweight models. Typically, lightweight networks do not perform as well as deeper networks. To solve this issue, knowledge distillation (KD), which aims to enable smaller (student) models to compete with larger (teacher) models in performance, has been introduced [14]. Due to its remarkable effectiveness in boosting the capability of lightweight models, KD has been widely used in various tasks, e.g., object detection [7, 39], semantic segmentation [24, 36], and natural language processing [30, 11].
@@
In summary, our main contributions are: 1) In KD, to account for the future benefits of adjusting instance temperature at the current stage, we formulate the instance temperature adjustment as a sequential decision-making task, and propose a novel method RLKD based on RL to handle this task. 2) To overcome the challenge of delayed rewards in our RLKD, we exploit a mechanism for instance reward calibration. Furthermore, we design a valid exploration strategy to promote the agent to learn valuable temperature adjustment policy with high efficiency. 3) Our RLKD can serve as a plug-and-play technique to boost the performance of KD algorithms. We validate its effectiveness on three benchmarks for image classification and semantic segmentation, all obtaining the state-of-the-art result.
@@ 3 Preliminary on Reinforcement Learning
$\vskip-1.99168ptG_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\vskip-1.13809pt$ (1)
@@ 4.2 Instance Reward Calibration
Here, $r^{b^{\prime}}_{n}$ represents the $n$-th corrected reward, and $n$ is the batch size. The variables $a$ and $b$ are weights, which we set to 1 and 0.5, respectively. The variable $G_{i}$ denotes the return at the $i$-th time step. Additionally, to ensure that the states $s^{b}$ recorded in the replay buffer match the corrected rewards, we devise a state updater $\mathcal{U}$ and update the states $s^{b}$ accordingly. The updated state $s^{b^{\prime}}$ is calculated as follows:
@@ 4.3 Efficient Exploration
Input: agent network $Q$, student network $f_{S}$, teacher network $f_{T}$, dataset $D$, batch size NN, reward corrector $\mathcal{C}$
1 for batch $d_{t+1}$ in $D$ do
2  for $i=0$ to N1N-1 do
3     Build state observation $s_{i}$ from $f_{S}(d_{t+1}^{i})$, $f_{T}(d_{t+1}^{i})$
4     Compute the action $a_{i}$ and values $V_{i}\leftarrow Q(s_{i})$
5     Use $a_{i}$ as instance temperature
6     Collect $s_{i}$, $a_{i}$, $V_{i}$ as a buffer to update $Q$ (Sec. 4.1)
7  end for
8 Update the student network $f_{S}$
9  Obtain the reward $r_{t+1}$ following Eq. 9
10  Calibrate the instance reward following Eq. 10
11  Update the reward $r_{t+1}$ in the replay buffer
12  Update the state $s_{t}$ following Eq. 12 (Sec. 4.2)
13  while not done do
14     Update $Q$
15  end while
16
17 end for
18Compute prediction entropy $H_{t}\leftarrow f_{S}(d_{t+1})$
19 Obtain high-quality samples following Eq. 16 (Sec. 4.3)
20 Implement the efficient exploration strategy
21 Update $Q$
@@ 5 Experiments
Tasks and datasets. Following [20], we conduct experiments on two tasks: image classification and object detection. For the image classification task, we carry out experiments on CIFAR-100 [18] and ImageNet [8]. For the object detection task, we conduct our experiments on MS-COCO [22]. CIFAR-100 is a prominent dataset for image classification, comprising $\times32$ pixel images across 100 different categories, with a training set of 50,000 images and a validation set of 10,000 images. ImageNet, another significant dataset for large-scale image classification, encompasses 1,000 categories with a training set of approximately 1.28 million images and a validation set of 50,000 images. MS-COCO is a famous dataset used for general object detection that includes 80 categories. It has a training set (train 2017) with 118,000 images and a validation set (val 2017) with 5,000 images.
@@ 5.1 Main results
Table 7: Ablation study of the efficient exploration ($E$) strategy.
@@ 5.2 Ablation studies
Table 8: Comparison of different high-quality training sample selection strategies. The teacher-student pairs corresponding to the second and third rows are respectively RN-56 & RN-20 and WRN-40-2 & WRN-16-2. $\mathcal{M}$” denotes the mix-up operation.